# Towards Understanding Clean Generalization and Robust Overfitting in Adversarial Training

### v3

@@ 1 Introduction
Second and most importantly, while some works (Tsipras et al., 2018; Zhang et al., 2019; Hassani & Javanmard, 2022) point out that achieving robustness may hurt clean test accuracy, in most of the cases, it is observed that drop of robust test accuracy is much higher than drop of clean test accuracy in adversarial training (Madry et al., 2017; Schmidt et al., 2018; Raghunathan et al., 2019) (see in Figure 1, where clean test accuracy is more than $80\%$ but robust test accuracy only attains nearly $50\%$). Namely, a weak version of benign overfitting (Zhang et al., 2017), which means that overparameterized deep neural networks can both fit random data powerfully and generalize well for unseen clean data, remains after adversarial training.
@@ 2 Additional Related Work
Adversarial Examples and Adversarial Training Method. Szegedy et al. (2013) first made an intriguing discovery: even state-of-the-art deep neural networks are vulnerable to adversarial examples. Then, Goodfellow et al. (2014) proposes FGSM to seek adversarial examples, and it also introduce the adversarial training framework to enhance the defence of models against adversarial perturbations.
@@ 6 Experiments
Model Size 1 4 8 12 16 1 2 5 10
Clean Test 11.35 11.35 11.35 95.06 94.85 82.56 84.92 85.83 86.05
Robust Test 11.35 11.35 11.35 77.96 83.43 43.39 43.74 46.25 50.08
Robust Train 11.70 11.70 11.70 99.30 99.50 64.19 79.82 97.37 99.57
In this section, we empirically verify our theoretical results in Section 4 and Section 5 by the numerical experiments in real-world image-recognition datasets and synthetic structured data, respectively.
@@ 6.1 Effect of Different Model Sizes
Experiment Settings. For the MNIST dataset, we consider a simple convolutional network, LeNet-5 (LeCun et al., 1998), and study how its performance changes as we increases the size of network (i.e. we expand the number of convolutional filters and the size of the fully connected layer to the size number multiple of the initial, where the size numbers are $1,4,8,12,16$). The original convolutional network has a convolutional layer with $1$ filters, followed by another convolutional layer with $2$ filters, and a fully connected hidden layer with 32 units. Convolutional layers are followed by $2\times 2$ max-pooling layers. For the CIFAR10 dataset, we apply WideResNet-34 (Zagoruyko & Komodakis, 2016) with different widen factors $1,2,5,10$. We use the standared projected gradient descent (PGD) (Madry et al., 2017) to train the network by adversarial training. We choose the classical $\ell_{\infty}$-perturbation with radius $0.3$ for MNIST and $8/255$ for CIFAR10. All models are trained for 100 epoches on MNIST and 200 epoches on CIFAR10 by using a single NVIDIA RTX 4090 GPU.
@@ 6.2 Synthetic Structured Data
Accuracy Train Test
Clean 100.0 98.5
Robust 100.0 17.5
@@ 7 Conclusion
To summarize, we study the puzzling clean generalization and robust overfitting (CGRO) phenomenon in adversarial training and present theoretical explanations: from the perspective of representation complexity, we prove that the CGRO classifier is efficient to achieve by leveraging robust memorization regarding the training data, while robust generalization requires excessively large model capacity in worst case, which may lead adversarial training to the CGRO regime; from the perspective of training dynamics, we propose a three-stage analysis technique to analyze the feature learning process of adversarial training under our structured-data framework, and it shows that two-layer neural network trained by adversarial training provably learns the partial true feature but memorizes the random noise from training data, which thereby causes the CGRO regime. On the empirical side, we confirm our theoretical findings above by real-world vision datasets and synthetic data simulation.

### v4

# On the Clean Generalization and Robust Overfitting in Adversarial Training from Two Theoretical Views: Representation Complexity and Training Dynamics

@@ 1 Introduction
Second and most importantly, while some works (Tsipras et al., 2018; Zhang et al., 2019; Hassani & Javanmard, 2022) point out that achieving robustness may hurt clean test accuracy, in most of the cases, it is observed that drop of robust test accuracy is much higher than drop of clean test accuracy in adversarial training (Madry et al., 2017; Schmidt et al., 2018; Raghunathan et al., 2019) (see in Figure 1, where clean test accuracy is more than $80\%$ but robust test accuracy only attains nearly $50\%$). Namely, a weak version of benign overfitting (Zhang et al., 2017), which means that overparameterized deep neural networks can both fit random data powerfully and generalize well for unseen clean data, remains after adversarial training. Therefore, it is natural to ask the following question:
@@ 2 Additional Related Work
Memorization in Adversarial Training. Dong et al. (2021); Xu et al. (2021) empirically and theoretically explore the memorization effect in adversarial training for promoting a deeper understanding of model capacity, convergence, generalization, and especially robust overfitting of the adversarially trained models. However, different from their works, the concept clean generalization and robust overfitting (CGRO) proposed in our paper focuses on both robust overfitting and high clean test accuracy, which means that there is surprisingly no clean memorization or clean overfitting.
@@ 6 Experiments
Model Size Factor 1 2 8 12 16 1 2 5 10
Clean Test Acc 11.35 11.35 11.35 95.06 94.85 82.56 84.92 85.83 86.05
Robust Test Acc 11.35 11.35 11.35 77.96 83.43 43.39 43.74 46.25 50.08
Robust Train Acc 11.70 11.70 11.70 99.3 99.5 64.19 79.82 97.37 99.57
In this section, we empirically verify our theoretical results in Section 4 and Section 5 by experiments in real-world image-recognition datasets and synthetic structured data.
@@ 6.1 Effect of Different Model Sizes
Experiment Settings. For the MNIST dataset, we consider a simple convolutional network, LeNet-5 (LeCun et al., 1998), and study how its performance changes as we increases the size of network (i.e. we expand the number of convolutional filters and the size of the fully connected layer to the size number multiple of the initial, where the size numbers are $1,4,8,12,16$). The original convolutional network has a convolutional layer with $1$ filters, followed by another convolutional layer with $2$ filters, and a fully connected hidden layer with 32 units. Convolutional layers are followed by $2\times 2$ max-pooling layers. For the CIFAR10 dataset, we apply WideResNet-34 (Zagoruyko & Komodakis, 2016) with different widen factors $1,2,5,10$. We use the standared projected gradient descent (PGD) (Madry et al., 2017) to train the network by adversarial training. We choose the classical $\ell_{\infty}$-perturbation with radius $0.3$ for MNIST and $8/255$ for CIFAR10. All models are trained for 100 epoches on MNIST and 200 epoches on CIFAR10.
@@ 6.2 Synthetic Structured Data
Train test
Clean Acc 100.0 98.5
Robust Acc 100.0 17.5
@@ 7 Conclusion
In this paper, we study the CGRO phenomenon in adversarial training and present theoretical explanations: from the perspective of representation complexity, we prove that the CGRO classifier is efficient to achieve by leveraging robust memorization regarding the training data, while robust generalization requires excessively large model capacity in worst case, which may lead adversarial training to the CGRO regime; from the perspective of training dynamics, we propose a three-stage analysis technique to analyze the feature learning process of adversarial training under our structured-data framework, and it shows that two-layer neural network trained by adversarial training provably learns the partial true feature but memorizes the random noise from training data, which thereby causes the CGRO regime. On the empirical side, we confirm our theoretical findings above by real-world vision datasets and synthetic data simulation.
Acknowledgements
Binghui Li is partially supported by the Elite Ph.D. Program in Applied Mathematics for PhD Candidates in Peking University. We thank anonymous reviewers for their valuable suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.